integration of- 1/(1-x2(1-x2)1/2)dx = ?

Grade:12

1 Answers

Suraj Prasad IIT Patna
9 years ago
assume x = sin\Theta, dx = cos\Theta d\Theta

so 1 – x^2 = cos^2\Theta

so integration becomes \intsec^2\Theta d\Theta = tan \Theta = tan (sin-1 x)